OpenStudy (anonymous):

Twice one number is 20 more than another number. One fourth the first number is 2 less than one fourth the second number. Find the numbers.

OpenStudy (mertsj):

x=first number y=2x-2 1/4x=1/4y-2

OpenStudy (anonymous):

does x=10 and y=18?

OpenStudy (mertsj):

Is twice 18 20 more than 1?

OpenStudy (anonymous):

geez srry that's what I got.!

OpenStudy (anonymous):

I don't get it...

OpenStudy (mertsj):

I actually left off a 0 y=2x-20

OpenStudy (anonymous):

am is supposed to use the substitution method to find the answer?

OpenStudy (anonymous):

am I???

OpenStudy (mertsj):

That's probably the easiest

OpenStudy (anonymous):

I keep getting a wrong answer.

OpenStudy (mertsj):

\[\frac{1}{4}x=\frac{1}{4}(2x-20)-2\] \[\frac{1}{4}x=\frac{1}{2}x-5-2\] \[-\frac{1}{4}x=-7\] \[x=28\] \[y=2x-20\] \[y=2(28)-20\] \[y=56-20\] \[y=36\]

OpenStudy (anonymous):

oh I did get the right answers. Thanks anyways!